# Repositioning, Ride-matching, and Abandonment in On-demand Ride-hailing Platforms: A Mean Field Game Approach

### v1

@@ 2.1. A Semi-Markov Decision Process
As the total mass is $m$, immediately, we obtain the conversation law of mass,
@@
(4) $\pi_{ik}(x)=\frac{x_{ik}}{\sum_{k\in A}x_{ik}}$
for state $i$ with a positive denominator, while $\pi_{ik}=0$ otherwise. In fact, the mass rate encapsulates the full information regarding drivers’ strategies. As we will demonstrate later in Proposition 2.2, the optimality of drivers’ strategies is derived in terms of the vector of mass rates.
Reward Structure. In this model, we assume a normalization of zero operating costs for vehicles and zero opportunity costs for drivers. We do not address the participation problem for drivers and assume that all drivers are willing to join the platform and adhere to the matching arrangements set by the platform. We assume that rewards are continuously generated at a compensatory rate of $c_{i}>0$ solely during the time when a driver is transporting a customer from the customer’s pick-up location to their destination. Consequently, with each trip originating from region $i$, a driver can, on average, earn $c_{i}t_{i}$. Given a stationary system, the average reward per driver per unit of time, denoted by $\Phi$, is a function of stationary mass rate vector $x$ given by
@@
This study focuses on the semi-Markov decision process with an average reward criterion where each driver aims to maximize their individual average reward over an infinite time horizon. Let $(\tilde{\pi}_{ik})_{i\in S,k\in A}$ be a reposition strategy for a representative driver. Let $\tilde{x}=(\tilde{x}_{ik})_{i\in S,k\in A}\in\mathbb{R}_{+}^{N^{2}}$ be the representative driver’s mass rate where for any $i$ and $k$, $\tilde{x}_{ik}$ represents the average number that this driver arrives in region $i$ and take action $k$ within one unit of time. Then $\tilde{\pi}$ and $\tilde{x}$ are related through a equation similar to (4),
@@
Given the system’s aggregated stationary mass rate $x$, the representative driver’s $\tilde{x}$ satisfies the conversation law and balance equation as follows,
(7a) $\sum_{i\in S}\tilde{x}_{ii}\big{(}w_{i}^{d}(x_{ii})+\tau_{i}(x_{ii})+t_{i}\big{)}+\sum_{i,j\in S\atop i\not=j}\tilde{x}_{ij}t_{ij}=1.$
(7b) $\sum_{j\in S}\tilde{x}_{ij}=\sum_{{j\in S},j\not=i}\tilde{x}_{ji}+\sum_{j\in S}\tilde{x}_{jj}q_{ji},\>\forall i\in S$
@@
The game $\mathcal{G}$ belongs to a mean field type. A mean field game (MFG) is a game-theoretical framework that models the strategic decision-making of a large number of interacting agents (Jovanovic and Rosenthal, 1988; Huang et al., 2006; Hopenhayn, 1992; Adlakha et al., 2015). In MFGs, each agent interacts with the ‘mean field’ (in our context, the vector $\mu$ of agent mass per state-action pair), which represents the average effect of the strategies taken by all the other agents. In our model, a key mean field term is the sojourn time $\big{(}T_{ij}(\cdot)\big{)}_{i,j\in S}$, which implicitly depends on the ‘mean field’ $\mu$ through the mass rate vector $x$. In this paper we make no assumptions that the reader is familiar with MFG theory, and our analysis is self-contained.
@@
Remark: We have not specified the size of each region in our setup. The region size will become relevant once we introduce the matching radius later in our analysis. Our methodology and analysis do not depend on the specific sizes of each individual region. It is more sensible to regard the mass we defined as mass per unit area (or mass density) and the rate as rate per unit area. Occasionally, we may omit the term “per unit area” when its context is apparent. To elucidate, we provide the physical meanings and suggested units for the primary variables and parameters in this paper, including those later introduced, in Table 1. These suggested units will be used in the numerical experiments and simulations conducted in this paper.
2.2. Stationary Equilibrium
Now we can give the definition of a stationary equilibrium of the game $\mathcal{G}$.
Definition 2.1.
  • (a)
  • (b)
Proposition 2.2.
(8) $\max\quad$ $\sum_{i\in S}x_{ii}c_{i}t_{i}$
(8c) over $x=(x_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
Proof.
Consider a mass rate $x^{\dagger}\in\mathbb{R}_{+}^{N^{2}}$, we will identify the conditions for $x^{\dagger}$ being an equilibrium mass rate. Note that we can compute the corresponding distribution $\mu^{\dagger}$ by (2). Hence, the first condition for $x^{\dagger}$ being an equilibrium mass rate, namely, Definition 2.1(a), is equivalent to that $x^{\dagger}$ satisfies (8a) and (8b) as per (3) and (1) respectively. Let $\tilde{x}=(\tilde{x}_{ik})_{i\in S,k\in A}\in\mathbb{R}_{+}^{N^{2}}$ be the representative driver’s mass rate. Since the representative driver’s strategy is encoded in $\tilde{x}$ through (6), he seeks to manipulate $\tilde{x}$ to maximize his average reward, i.e. $\sum_{i\in S}\tilde{x}_{ii}c_{i}t_{i}$. A feasible $\tilde{x}$ must satisfies (7b) in a stationary system. Moreover, as the sojourn time is determined by the aggregate mass rate $x^{\dagger}$, similar to (7a), a feasible $\tilde{x}$ must satisfy that
Thus, the representative driver’s optimization problem is,
(9) $\max\quad$ $\sum_{i\in S}\tilde{x}_{ii}c_{i}t_{i}$
(9c) over $\tilde{x}=(\tilde{x}_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
Theorem 2.3.
Suppose Assumption 1 is true, an $x^{\dagger}\in\mathbb{R}_{+}^{N^{2}}$ is an equilibrium mass rate if and only if $x^{\dagger}$ satisfies the KKT conditions of the optimization problem:
over $x=(x_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
Proof.
(11) $m\log\bigg{(}\sum_{i}x_{ii}c_{i}t_{i}\bigg{)}.$
Thus, $x^{\dagger}$ satisfies the first order conditions:
$+\sum_{i}x^{\dagger}_{ii}\sum_{j\not=i}(\nu_{i}-\nu_{j})q_{ij}+\sum_{j\not=i}(\nu_{i}-\nu_{j})x^{\dagger}_{ij}=0$
Apply the conservation law (8a) and balance equation (8b), we obtain $m(1-\lambda)=0$, so $\lambda=1$. Therefore, the first order conditions can be rewritten as
(13a) $\frac{mc_{i}t_{i}}{\sum_{i}x^{\dagger}_{ii}c_{i}t_{i}}-t_{i}-\tau_{i}(x^{\dagger}_{ii})-w_{i}^{d}(x^{\dagger}_{ii})+\sum_{j\not=i}(\nu_{i}-\nu_{j})q_{ij}\leq 0,\text{ with equality if }x^{\dagger}_{ii}>0,\ \forall i$
(13b) $-t_{ij}+\nu_{i}-\nu_{j}\leq 0,\text{ with equality if }x^{\dagger}_{ij}>0,\ \forall i,j,i\not=j$
Conversely, if $x^{\dagger}$ satisfies the KKT conditions of the optimization problem (10), i.e., (13a) and (13b), it also satisfies the first order conditions (12a) and (12b) for optimization problem (8) with $\lambda=1$ and objective function replaced by (11) . By multiplying both sides of (13a) and (13b) by the respective $x^{\dagger}_{ij}$ for all $i$, $j$, then summing them up, we obtain the conservation law (8a). Since $x^{\dagger}$ also satisfies the balance equation (8b), $x^{\dagger}$ is feasible and hence optimal for optimization problem (8). Thus, $x^{\dagger}$ is an equilibrium. ∎
Corollary 2.4.
Proof.
It suffices to show that there exists at least one optimal solution to (10). This is because optimization problem (10) has linear constraints and hence, any optimal solution must adhere to the Karush–Kuhn–Tucker conditions. Then the existence of an equilibrium follows from Theorem 2.3. The current domain for variable $x$ in (10), $\mathbb{R}_{+}^{N^{2}}$, is open and unbounded. We need to show that the optimal value of (10) can be obtained in a closed and bounded (compact) subset of $\mathbb{R}_{+}^{N^{2}}$.
First, it should be noted that the optimal value of (10) must be obtained away from the zero vector, which is not a feasible solution. This is because as $||x||\rightarrow 0$, the objective value tends towards negative infinity. Second, as per Assumption 1, the mass rate $x_{ii}$ has an strict upper bound $b_{i}$ for any $i$. Third, for any $i\not=j$, as $x_{ij}$ goes to infinity, the objective function in (10) goes to negative infinity since $t_{ij}>0$. Consequently, the optimal value of (10) must be obtained when any $x_{ij}$ with $i\not=j$ is finite. Furthermore, the optimal value must be obtained when $x_{ii}$ is strictly less than $b_{i}$ for any $i$ since the derivative of the objective function in (10) with respect to $x_{ii}$,
Corollary 2.5.
Suppose i) Assumption 1 holds, and ii) the total sojourn time of the driver, $T_{i}(x_{ii})$, is a strictly increasing function of $x_{ii}$ for every $i\in S$. Then, a unique stationary equilibrium exists for the game $\mathcal{G}$.
On the other hand, if $T_{i}(x_{ii})$ is not an increasing function of $x_{ii}$, multiple equilibria can arise. In fact, under most dispatch algorithms, the pick-up time is not a monotonic function of $x_{ii}$. This non-monotonicity leads to the “Wild Goose Chase” equilibrium, where drivers may be dispatched to pick up customers located far away. This issue has also been investigated in prior work (Castillo et al., 2023). We will provide a more detailed mathematical analysis of WGC to investigate the impact of dispatch algorithms. We propose a novel dispatch algorithm designed to ensure that $T_{i}(x_{ii})$ is an increasing function of $x_{ii}$. This guarantees a unique equilibrium and eliminates WGC equilibrium.
3. Ride-Matching Algorithm under Abandonment Behavior
In this section, we present a comprehensive analysis of the platform’s ride-matching algorithms and customer abandonment behavior, examining their influence on the equilibrium and overall performance of the ride-hailing system. We show that the widely used nearest-neighbor dispatch, either without a matching radius or with an arbitrarily chosen one, can result in “Wild Goose Chase” equilibrium and the emergence of multiple equilibria. This underscores the importance of optimizing the design of the matching radius, a topic we explore in the next section.
@@ 3.1. Customer: Abandonment Behavior
Ride-hailing platforms typically permit customers to cancel the service if they consider their wait time to be excessively long. Our model, similar to (Yang et al., 2020; Wang et al., 2022), assumes that each customer has a maximum willing waiting time. If a customer is not matched with a driver within this time threshold, they will exit the system. Additionally, we assume that once a match is made, customers are committed to the ride and cannot abandon the service.
Suppose that each waiting customer in region $i$ has an exponential clock with parameter $\theta_{i}$ that if it expires prior to the customer being matched with a driver, this customer abandons the system. In the spirit of mean-field approximation, if at any time the mass density of waiting customer is $\mu_{i}^{c}$, these will generate a rate of abandonment $\theta_{i}\mu_{i}^{c}$, where
(14) $\theta_{i}\mu_{i}^{c}+x_{ii}=b_{i}$
i.e., abandonment rate + service rate (driver supply rate) = arrival rate. This equation establishes a relationship between the unknown variables $x_{ii}$, $\mu_{i}^{c}$ and known parameters $b_{i}$, $\theta_{i}$ using mean-field approximation.
@@ 3.2. Platform: Matching Policy
We assume the platform adopt a notable dispatch algorithm: nearest-neighbor dispatch (NN), which is simple and near-optimal (Besbes et al., 2022). Recent work propose using a matching radius to prevent drivers from being dispatched to pick up customers who are too far away, thereby improve the system performance (Castillo et al., 2023; Yang et al., 2020). Our observation is that only one matching radius is lack of flexibility. A key innovation of this study is the introduction of a two-matching-radius nearest-neighbor dispatch algorithm:
  • When a new customer arrives (as in Fig. 1.(a) ), the platform assigns him to the closest available driver within the first matching radius ($R_{c}$).
  • When a driver becomes available (as in Fig. 1.(b) ), the platform assigns him to the closest waiting customer within a second matching radius ($R_{d}$), which may differ from the first matching radius.
@@
3.3. Formulas of Waiting Time and Pick-up Time

### v2

@@ 2.1. A Semi-Markov Decision Process
As the total mass is $m$, immediately, we obtain the conservation law of mass,
@@
(4) $\pi_{ik}(x)=\frac{x_{ik}}{\sum_{k\in A}x_{ik}}\,,$
for state $i$ with a positive denominator, while $\pi_{ik}=0$ otherwise. In fact, the mass rates encapsulate the full information regarding drivers’ strategies. As we will demonstrate later in Proposition 2.2, the optimality of drivers’ strategies is derived in terms of the vector of mass rates.
Reward Structure. In this model, we assume [^2] We can extend the cost model and our analysis still holds. a normalization of zero operating costs for vehicles and zero opportunity costs for drivers. We do not address the participation problem for drivers and assume that all drivers are willing to join the platform and adhere to the matching arrangements set by the platform. We assume that rewards are continuously generated at a compensatory rate of $c_{i}>0$ solely during the time when a driver is transporting a customer from the customer’s pick-up location to their destination. Consequently, with each trip originating from region $i$, a driver can, on average, earn $c_{i}t_{i}$. Given a stationary system, the average reward per driver per unit of time, denoted by $\Phi$, is a function of the stationary mass rate vector $x$ given by
@@
This study focuses on the semi-Markov decision process with an average reward criterion where each driver aims to maximize their individual average reward over an infinite time horizon. Let $(\tilde{\pi}_{ik})_{i\in S,k\in A}$ be a reposition strategy for a representative driver. Let $\tilde{x}=(\tilde{x}_{ik})_{i\in S,k\in A}\in\mathbb{R}_{+}^{N^{2}}$ be the representative driver’s mass rate where for any $i$ and $k$, $\tilde{x}_{ik}$ represents the average number of times that this driver arrives in region $i$ and takes action $k$ within one unit of time. Then $\tilde{\pi}$ and $\tilde{x}$ are related through an equation similar to (4),
@@
Given the system’s aggregated stationary mass rate $x$, the representative driver’s $\tilde{x}$ satisfies the conservation law and balance equation as follows,
(7a) $\sum_{i\in S}\tilde{x}_{ii}\big{(}w_{i}^{d}(x_{ii})+\tau_{i}(x_{ii})+t_{i}\big{)}+\sum_{i,j\in S\atop i\not=j}\tilde{x}_{ij}t_{ij}=1\,,$
(7b) $\sum_{j\in S}\tilde{x}_{ij}=\sum_{{j\in S},j\not=i}\tilde{x}_{ji}+\sum_{j\in S}\tilde{x}_{jj}q_{ji},\>\forall i\in S\,.$
@@
The game $\mathcal{G}$ belongs to a mean field type. A mean field game (MFG) is a game-theoretical framework that models the strategic decision-making of a large number of interacting agents. In MFGs, each agent interacts with the ‘mean field’ (in our context, the vector $\mu$ of agent mass per state-action pair), which represents the average effect of the strategies taken by all the other agents. In our model, a key mean field term is the sojourn time $\big{(}T_{ij}(\cdot)\big{)}_{i,j\in S}$, which implicitly depends on the mean field $\mu$ through the mass rate vector $x$. In this paper, we make no assumptions that the reader is familiar with MFG theory, and our analysis is self-contained.
2.2. Stationary Equilibrium
We give the definition of a stationary equilibrium of the game $\mathcal{G}$.
Definition 2.1.
  • (a)
  • (b)
Proposition 2.2.
(8) $\max\quad$ $\sum_{i\in S}x_{ii}c_{i}t_{i}$
(8c) over $x=(x_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
Proof.
Consider a mass rate $x^{\dagger}\in\mathbb{R}_{+}^{N^{2}}$, we will identify the conditions for $x^{\dagger}$ being an equilibrium mass rate. Note that we can compute the corresponding distribution $\mu^{\dagger}$ by (2). Hence, the first condition for $x^{\dagger}$ being an equilibrium mass rate, namely, Definition 2.1(a), is equivalent to that $x^{\dagger}$ satisfies (8a) and (8b) as per (3) and (1) respectively. Let $\tilde{x}=(\tilde{x}_{ik})_{i\in S,k\in A}\in\mathbb{R}_{+}^{N^{2}}$ be the representative driver’s mass rate. Since the representative driver’s strategy is encoded in $\tilde{x}$ through (6), he seeks to manipulate $\tilde{x}$ to maximize his average reward, i.e. $\sum_{i\in S}\tilde{x}_{ii}c_{i}t_{i}$. A feasible $\tilde{x}$ must satisfy (7b) in a stationary system. Moreover, as the sojourn time is determined by the aggregate mass rate $x^{\dagger}$, similar to (7a), a feasible $\tilde{x}$ must satisfy that
Thus, the representative driver’s optimization problem is,
(9) $\max\quad$ $\sum_{i\in S}\tilde{x}_{ii}c_{i}t_{i}$
(9c) over $\tilde{x}=(\tilde{x}_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
Theorem 2.3.
Suppose Assumption 1 is true, an $x^{\dagger}\in\mathbb{R}_{+}^{N^{2}}$ is an equilibrium mass rate of game $\mathcal{G}$ if and only if $x^{\dagger}$ satisfies the KKT conditions of the optimization problem:
over $x=(x_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
Proof.
(11) $m\log\bigg{(}\sum_{i}x_{ii}c_{i}t_{i}\bigg{)}.$
Thus, $x^{\dagger}$ satisfies the first order conditions:
$+\sum_{i}x^{\dagger}_{ii}\sum_{j\not=i}(\nu_{i}-\nu_{j})q_{ij}+\sum_{j\not=i}(\nu_{i}-\nu_{j})x^{\dagger}_{ij}=0\,.$
Applying the conservation law (8a) and balance equation (8b), we obtain $m(1-\lambda)=0$, so $\lambda=1$. Therefore, the first order conditions can be rewritten as
(13a) $\frac{mc_{i}t_{i}}{\sum_{i}x^{\dagger}_{ii}c_{i}t_{i}}-t_{i}-\tau_{i}(x^{\dagger}_{ii})-w_{i}^{d}(x^{\dagger}_{ii})+\sum_{j\not=i}(\nu_{i}-\nu_{j})q_{ij}\leq 0,\text{ with equality if }x^{\dagger}_{ii}>0,\ \forall i\,,$
(13b) $-t_{ij}+\nu_{i}-\nu_{j}\leq 0,\text{ with equality if }x^{\dagger}_{ij}>0,\ \forall i,j,i\not=j\,,$
Conversely, if $x^{\dagger}$ satisfies the KKT conditions of the optimization problem (10), i.e., (13a) and (13b), it also satisfies the first order conditions (12a) and (12b) for the optimization problem (8) with $\lambda=1$ and objective function replaced by (11) . By multiplying both sides of (13a) and (13b) by the respective $x^{\dagger}_{ij}$ for all $i$, $j$, then summing them up, we obtain the conservation law (8a). Since $x^{\dagger}$ also satisfies the balance equation (8b), $x^{\dagger}$ is feasible and hence optimal for the optimization problem (8). Thus, $x^{\dagger}$ is an equilibrium. ∎
Corollary 2.4.
Proof.
By Theorem 2.3, we only need to show that there exist KKT points. Because optimal solutions to (10) are KKT points, it suffices to show there exist such solutions. Our function is continuous and if maximized over a compact region, it would have an optimal solution. But since our domain is not compact, we need to show that optimal values can be obtained in a compact subset of the function domain. The current domain for variable $x$ in (10), $\mathbb{R}_{+}^{N^{2}}$, is open and unbounded. We need to show that the optimal value of (10) can be obtained in a closed and bounded (compact) subset of $\mathbb{R}_{+}^{N^{2}}$.
First, it should be noted that the optimal value of (10) must be obtained away from the zero vector, which is not a feasible solution. This is because as $||x||\rightarrow 0$, the objective value tends towards negative infinity. Second, as per Assumption 1, the mass rate $x_{ii}$ has a strict upper bound $b_{i}$ for any $i$. Third, for any $i\not=j$, as $x_{ij}$ goes to infinity, the objective function in (10) goes to negative infinity since $t_{ij}>0$. Consequently, the optimal value of (10) must be obtained when any $x_{ij}$ with $i\not=j$ is finite. Furthermore, the optimal value must be obtained when $x_{ii}$ is strictly less than $b_{i}$ for any $i$ since given part $(2)$ in Assumption 1, the derivative of the objective function in (10) with respect to $x_{ii}$
Corollary 2.5.
Suppose i) Assumption 1 holds, and ii) the total sojourn time of a driver, $T_{i}(x_{ii})$, is a strictly increasing function of $x_{ii}$ for every $i\in S$. Then, a unique stationary equilibrium exists for the game $\mathcal{G}$.
On the other hand, if $T_{i}(x_{ii})$ is not an increasing function of $x_{ii}$, multiple equilibria can arise. In fact, under most dispatch algorithms, the pick-up time is not a monotonic function of $x_{ii}$. In the case of multiple equilibria, the one with the lowest $x_{ii}$ is not efficient since drivers may be dispatched to pick up customers located far away and many customer abandon the system due to longer waiting. In Section 3.4 we analyze the impact of driver dispatch algorithms, and propose a novel dispatch algorithm designed to ensure that $T_{i}(x_{ii})$ is an increasing function of $x_{ii}$. This guarantees a unique equilibrium and eliminates inefficient equilibria.
3. Ride-Matching Algorithms and Analysis of Equilibria
In this section, we present a comprehensive analysis of the platform’s ride-matching algorithms and customer abandonment behavior, examining their influence on the equilibria and the resulting overall performance of the ride-hailing system. We begin by introducing the customer abandonment behavior model in Section 3.1. Next, we present our design of the platform’s matching algorithm in Section 3.2. Section 3.3 derives formulas for driver waiting and pick-up times as functions of the driver supply rate, a prerequisite (Assumption 1) for applying Theorem 2.3 to identify equilibria as KKT points of an optimization problem. Building on these results, Section 3.4 demonstrates that the widely adopted nearest-neighbor dispatch approach–whether implemented without a matching radius or with an arbitrarily chosen one–can lead to multiple equilibria, some of which are characterized by lengthy pick-up times. This finding emphasizes the critical need to optimize the design of the matching radius, a topic we delve into in the following section.
@@ 3. Ride-Matching Algorithms and Analysis of Equilibria
Remark on Region Size and Units. In Section 2, we did not specify the size of each physical region because the methodology and analysis presented there are independent of the specific dimensions of individual regions. However, the size of each region becomes relevant when we introduce the matching radius in this section. Performance metrics of the ride-hailing system—such as waiting times and pick-up times—are sensitive to both the size of regions and the mass per unit area within them. To reflect this, we now redefine the mass and rate introduced in Section 2 as mass per unit area (or mass density) and rate per unit area, respectively. For brevity, we may occasionally omit the term ”per unit area” when the context is clear. For clarity, Table 1 summarizes the physical meanings and suggested units for the primary variables and parameters used in this paper, including those introduced later. These units will be employed in the numerical experiments and simulations conducted throughout the study.
@@ 3.1. Customer: Abandonment Behavior
We introduce the the customer abandonment behavior model in this section. Ride-hailing platforms typically permit customers to cancel the service if they consider their waiting time to be excessively long. Our model, similar to (Yang et al., 2020; Wang et al., 2022), assumes that each customer has a (random) maximum willing waiting time. If a customer is not matched with a driver within this time limit, they will exit the system. Additionally, we assume that once a match is made, customers are committed to the ride and cannot abandon the service.
In our equilibrium analysis we assume that each waiting customer in region $i$ has an exponential alarm clock with parameter $\theta_{i}$, that if it expires prior to the customer being matched with a driver, this customer abandons the system. Thus, $\theta_{i}$ represents the impatience level of customers in region $i$: a higher value of $\theta_{i}$ indicates greater impatience among customers. In the spirit of mean-field approximation, if the mass density of waiting customers is $\mu_{i}^{c}$, these will generate a rate of abandonment $\theta_{i}\mu_{i}^{c}$, where
(14) $\theta_{i}\mu_{i}^{c}+x_{ii}=b_{i}\,,$
i.e., customer abandonment rate + service rate (driver supply rate) = customer arrival rate in an equilibirium state. This equation establishes a relationship between the unknown variables $x_{ii}$, $\mu_{i}^{c}$ and known parameters $b_{i}$, $\theta_{i}$ using mean-field approximation.
@@ 3.2. Platform: Matching Policy
We present our design for the platform’s matching algorithm in this section. We assume the platform adopts ‘nearest-neighbor dispatch’ (NN), which is simple and near-optimal (Besbes et al., 2022). Recent work proposes using a matching radius to prevent drivers from being dispatched to pick up customers who are too far away, thereby improving the system performance (Castillo et al., 2023; Yang et al., 2020). Our observation is that only one matching radius lacks flexibility. A key innovation of this study is the introduction of a two-matching-radius nearest-neighbor dispatch algorithm:
  • When a new customer arrives (as in Fig. 1.(a) ), the platform assigns him to the closest available driver within the first matching radius $R_{c}$.
  • When a driver becomes available (as in Fig. 1.(b) ), the platform assigns him to the closest waiting customer within a second matching radius $R_{d}$, which may differ from the first matching radius.
@@
3.3. Formulas for Waiting Times and Pick-up Times